Grade 12th passElectrostatics

Two identical balls A&B are moving with same velocity.If velocity of A is reduced to half and half and of `B` to zero then the rise in temperatures of A to that of `B` is

Rehana
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Let m be the mass and v be the initial velocity of the balls. Let s be their specific heat capacity.
Energy dissipated for ball A = 
= Initial KE - Final KE
= 1/2 mv^2 - 1/2 m (v/2)^2
= 3/8 mv^2
 
Now m*s* \DeltatA= 3/8 mv^2
Hence
\Delta tA = 3 v^2/ 8s
 
Energy dissipated for ball B = 
= Initial KE - Final KE
= 1/2 mv^2 - 0
= 1/2 mv^2
 
Now 
m*s* \Delta tB = 1/2 mv^2
Hence
 \DeltatB= v^2/ 2s
 
Hence
\DeltatA/\DeltatB = 3/4